The balanced equation for the decomposition of TNT, C₇ H₅(NO₂)₃, is


2 C₇ H₅(NO₂)₃(s) → 7 C(s)+7 CO(g)+3 N₂(g)+5 H₂ O(l)


How many moles of each product would form if 8.32 mol of C₇ H₅(NO₂)₃ react?

_______ mol C

_______ mol CO

_______ mol N₂

_______ mol H₂ O
Answer #1

From the equation, moles of C = 7 / 2 * moles of C7H5(NO2)3

= 7 / 2 * 8.32 = 29.12 moles

Moles of CO = 7 / 2 * moles of C7H5(NO2)3 = 7 / 2 * 8.32 = 29.12 moles

Moles of N2 = 3 / 2 * moles of C7H5(NO2)3 = 3 / 2 * 8.32 = 12.48 moles

Moles of H2O = 5 / 2 * moles of C7H5(NO2)3 = 2.5 * 8.32 = 20.8 moles